OpenStudy (anonymous):

Find the derivative of f(x) = -(9/x) at x = -8. How would I solve this?

OpenStudy (anonymous):

\[-(\frac{ 9 }{ x })=-9x^{-1}\] Then use the power rule \[\frac{ d }{ dx }x^n=nx^{n-1}\]

OpenStudy (usukidoll):

either use quotient rule or the power rule

OpenStudy (anonymous):

Thanks guys. I think I get it

OpenStudy (usukidoll):

alright :)

OpenStudy (anonymous):

Can you tell me if I have the correct answer?

OpenStudy (usukidoll):

sure.

OpenStudy (anonymous):

8/9?

OpenStudy (usukidoll):

let me check.. I"m going to use power rule as @peachpi did

OpenStudy (usukidoll):